Upon arrival I found ITF in Commissioning Mode and in SINGLE_BOUNCE_WI.
At 13:39 UTC I set Upgrading Mode.
At 14:00 UTC I started the planned operations before the break of the vacuum of SDB1:
- ITF in DOWN, relock IR cavities, DOWN again.
- Verified that DET_MAIN metatron node status was in SHUTTER CLOSED, DET metatron node set in PAUSE.
- SDB2 suspension control opened via VPM process SDB2_SBE > Open Position Loop.
- Angular control loops of SDB1 opened via VPM process SDB1_LC > Marionette > Open Angular Loop.
- Suspension control loops of SDB1 (ID) opened.
- OMC temperature control switched off via VPM process SDB1_OMC > Temp Stab cmd > Disable (-1)
During the operations, the SQB1_SBE loop opened. According with SQZ expert, I closed it.
15:30 UTC ITF relocked in LOCKED_ARMS_IR. Still there.
Vacuum
SDB1 venting in progress (Vacuum team).
Susp
SR measurements done from 14:30 to 15:00 UTC (Boschi).
SQZ
Under request of DeLaurentis I recovered the vertical position of SQB2 via stepping motors.
I tried also to recover the FCIM loops but the horizontal position was far from the correct working point. According with DeLaurentis we decided to ask Bertolini to check on it tomorrow morning.